TT b) x=sec?t-1; ly= = tant at t = 1=- 4determine the equation of a tangent line to the curve
Answer #1

(6) S&= (a= see?t-1; y tant at 7 =- 21 Ať to- 4 X 송 seca(-7) -1 = 2-1 = 1 Y = tan (-7)=-4 Now, dac de - 2 seet. Seit tant = 2